Report 1.

Make a table to show and compare the measured pH using the pH meter and the paper for each solution to the calculated pH.

#

Solution

pH

(paper)

pH

(pH meter)

pH

(Calculated)

1

0.1 M HCl

1.0

1.115

2

0.01 M HCl

4.0

2.925

3

0.001 M HCl

3.0

2.942

4

0.1 M NaCl

5.0

5.809

5

0.1 M HC2H3O2

3.0

2.932

6

0.1 M NaOH

12.0

12.793

7

deionized water

6.0

6.664

8

Tap water

7.0

7.722

9

0.1 M NH3

9.0

10.376

10

0.1 M NH4Cl

6.0

6.275

11

0.1 M (NH4)2C2O4

8.0

7.954

12

0.1 M NaC2H3O2

7.0

7.965

13

0.1 M NaH2PO4

4.0

4.421

14

0.1 M Na2HPO4

9.0

8.952

15

0.1 M Na3PO4

12.0

12.259

16

0.1 M H3PO4

1.0

1.714

  

  1. Calculate the pH for each of the above solutions.    

Report

       Prepare a table with volume of NaOH or HCl added and the resulting pH of the water and the buffer solution.  

Water

Buffer solution

Beaker 1

Beaker 2

Beaker 3

Beaker 4

Total Volume of

1 M NaOH

added

pH

Total Volume of

1 M HCl

added

pH

Total Volume of

1 M NaOH

added

pH

Total Volume of

1 M HCl

added

pH

0 mL

4.0

0 mL

4.0

0 mL

4.740

0 mL

4.674

1 drop

8.0

1 drop

3.509

1 mL

4.714

1 mL

4.527

2 drops

7.0

2 drops

3.371

2 mL

4.715

2 mL

4.302

6 drops

9.0

6 drops

2.923

6 mL

4.735

6 mL

3.806

10 drops

9.0

10 drops

2.694

10 mL

4.764

10 mL

2.585

20 drops

8.0

20 drops

2.406

20 mL

4.886

20 mL

0.979

2 mL =
(40 drops)

11.0

2 mL

2.139

30 mL

5.073

30 mL

0.819

  1. Calculate the concentration of the sodium acetate and acetic acid buffer solution,
  2. Calculate the expected pH of the buffer solution. Compare this to the measured pH

  1. Create a table with the mass of KHP and volume of NaOH solution used in each titration. Show the moles of KHP, moles of NaOH and molarity of NaOH.

  2. Show how to calculate the moles of KHP, moles of NaOH and molarity of the NaOH solution using your values. . Use at least 6 trials or the complete class data to determine the average and standard deviation of the molarity.

Answer the following question:  

  1. If we know the mass of NaOH used to make the solution, why did we use the primary standard to determine the concentration of the NaOH solution?
Answer #1

two questions/Reports , with so many sub-parts (calculation ) as per guideline only 1st report question will be answered (completely ) :

pH calculation for each solution :

1. 0.1 M HCl

pH = - (log[H+ ]) = - (log[0.1 ]) = 1.0

2. 0.01 M HCl

pH = - (log[H+ ]) = - (log[0.01 ]) = 2.0'

3. 0.001 M HCl

pH = - (log[H+ ]) = - (log[0.001 ]) = 3.0

4. 0.1 M NaCl (it is salt of strong base and strong acid , thus solution will be neutral , as salt hydrolysis will give equal amount of H+ and OH- )

pH = 7.0

5. 0.1 M CH3COOH (weak acid ) ka = 1.8*10-5

for weak acid [H+ ] = (ka*(HA))^1/2 = (1.8*10-5*(0.1))^1/2 = 1.34*10-3 M

pH = - (log[H+ ]) = 2.87

6. 0.1 M NaOH

[OH- ] = 0.1 M

pH = 14 -pOH = 14 - (-log[0.1 ] = 13

7. De-ionized water kw =  [H+ ][OH- ] = 10-14

[H+ ]= 10-7

pH = - (log[H+ ]) =7.00

8. tap water may contain impurities, which may make it acidic or basic : so, it cannot be calculated without any other reference.

9. 0.1 M NH3 (weak base ) kb = 1.8*10-5

[OH- ] = (kb*(B))^1/2 = (1.8*10-5*(0.1))^1/2 = 1.34*10-3 M

pH = 14 -pOH = 14 - (-log[1.34*10-3 ] = 11.13

10. 0.1 M NH4Cl (salt of weak base and strong acid : solution will be acidic)

NH4+ (aq)+ 2H2O (l) <===>NH4OH (aq)+ H3O+ (aq)

kb = 1.8*10-5

[H+ ]= ((kh (NH4Cl ))^1/2 = ((kw/kb* (NH4Cl ))^1/2 = ((5.6*10-10 (0.1 ))^1/2 = 7.48*10-6

pH = - (log[H+ ]) =5.13

11. 0.1 M (NH4)2C2O4 (salt of weak base and weak acid : pH will depend on Ka and Kb values : for oxalic acid we will consider ka2 )

kb = 1.8*10-5 ; Ka2 = 1.5*10-4

kh = (kw / kb*ka2) = 3.7*10-6

[H+ ]= (kh)^1/2 ka2 = (3.7*10-6)^1/2*1.5*10-4 = 2.89*10-7

pH = - (log[H+ ]) = 6.54

12. 0.1 M  CH3COONa (Salt of weak acid and strong base : solution will be basic)

CH3COO- (aq)+H2O (l) <===>CH3COOH (aq)+ OH- (aq)

ka = 1.8*10-5

[H+ ]= = ((kw* ka /(CH3COONa ))^1/2 = (( 1.8*10-19 /(0.1 ))^1/2 =1.34*10-9

pH = - (log[H+ ]) =8.87

13. 0. 1 M NaH2PO4  

(Salt of weak cojugated cation and amphoteric anion : k1 (7.5*10-3) ,k2 (6.2*10-8),k3 (10-12)for H3PO4 will be used )

pH = 1/2 (pk1 + pk2) = 4.66

14.

0. 1 M Na2HPO4

(Salt of weak cojugated cation and amphoteric anion : k1,k2 (6.2*10-8),k3 (10-12)for H3PO4 will be used )

pH = 1/2 (pk2 + pk3) = 9.6

15.

0. 1 M Na3PO4  

(Salt of weak acid and strong base : solution will be basic)

PO43-(aq)+H2O (l) <===> HPO42-(aq)+ OH- (aq)

ka3 = 10-12

[H+ ]= = ((kw* ka /(CH3COONa ))^1/2 = (( 10-26 /(0.1 ))^1/2 =3.16*10-13

pH = - (log[H+ ]) =12.5

16

0. 1 M H3PO4  ( weak acid : ka1 = 7.5*10-3))

for weak acid [H+ ] = (ka*(HA))^1/2 = (7.5*10-3)*(0.1))^1/2 = 2.74*10-2 M

pH = - (log[H+ ]) = 1.56